1. The demand for petrol rises from 500 to 600 Barrels when the price of a particular scooter is reduced from Birr. 25000 to Birr.22000. Find out the cross elasticity of demand for the two. What is the nature of their relationship?
A company has the following demand equation
Q= 1000–3000P+10A
Q = Quantity demanded
P = Product Price
A = Advertisement expenditure
Assume that P = 3 and A = 2000
2. Suppose the firm drops the price to Birr. 2.50 would this bebeneficial.
3. Suppose the firm raises the price to Birr. 4.00 While increasingits advertisement expenditure by 100 would this be beneficial? Explain
1
Expert's answer
2020-09-03T13:49:46-0400

1. If the demand for petrol rises from 500 to 600 Barrels when the price of a particular scooter is reduced from Birr. 25000 to Birr. 22000, then the cross elasticity of demand for the two is:

"Ed = \frac{600 - 500}{22000 - 25000}\times\frac{22000 + 25000}{600 + 500} = \frac{100}{-3000}\times\frac{47000}{1100} = -1.42."

So, these products are complements.

A company has the following demand equation

Q= 1000–3000P+10A

Q = Quantity demanded

P = Product Price

A = Advertisement expenditure

If P = 3 and A = 2000, then "Q1 = 1000 - 3000*3 + 10*2000 = 12000."

"TR1 = P1*Q1 = 3*12000 = 36000."

2. If the firm drops the price to Birr. 2.50, then "Q2=1000\u22123000\u22172.5+10\u22172000=13500."

"TR2 = P2*Q2 = 2.5*13500 = 33750."

So, this decrease would not be beneficial, because the demand is inelastic.

3. If the firm raises the price to Birr. 4.00 while increasing its advertisement expenditure by 100, it would be beneficial.